# M4FC: a Multimodal, Multilingual, Multicultural, Multitask Real-World Fact-Checking Dataset

### v1

@@ 5 Conclusion
We introduce M4FC, a real-world dataset for multimodal fact-checking. The images in M4FC have a broad geographic coverage, representing diverse cultures and world events. The dataset supports 10 claim languages and spans six tasks, including two new ones, broadening the scope of multimodal AFC. Our baseline results show that M4FC is a challenging dataset for MLLMs and SOTA AFC methods. Furthermore, we showed that intermediate tasks can have a high impact on downstream verdict prediction performance.
@@
We identify four limitations to M4FC.
First, M4FC does not cover videos and audio. Videos have recently become the most prevalent misinformation modality verified by fact-checkers (Dufour et al., 2024), due to the rise of social media platforms like TikTok. However, it is challenging to obtain the videos used in real-world multimodal claims. The fact-checking articles usually only include a screenshot, and the videos themselves are often removed from the social media platforms where they appeared. Implementing a data collection pipeline to create a real-world dataset like M4FC, but with videos instead of images, is an important direction for future work.
Second, despite the creation of scraping scripts customized for each fact-checking organization, a lot of manual verification was needed. For example, many image URLs had to be corrected. We also did not cover fact-checking web pages that use JavaScript. This limits the possibilities to automatically and easily scale the dataset. Future work should consider better automated ways to scrape the raw content of fact-checking web pages.
@@
Fourth, the balanced and multilingual verdict prediction settings may suffer from biases. To avoid relying on news articles to source true claims, we generated additional true claims with GPT4o based on the context items of the image. This introduces some biases: the generated true claims are all formulated in a similar way and they are often more detailed than false claims, e.g., they contain specific dates. A trained verdict prediction classifier might exploit these biases. Hence, we do not recommend training on the balanced setting for verdict prediction, but rather using only its test set, after training on the imbalanced train set or any other dataset. Furthermore, it is not recommended to train a classifier on the multilingual and imbalanced setting, as most true claims are contained within two languages, English and Portuguese, while other languages only contain false claims, like Arabic. This verdict imbalance by language might also be exploited as a shortcut by a trained classifier.
@@ Acknowledgments
Appendix A Labeling prompts
Table 7 presents the prompt instructions provided to GPT4o to extract task labels and metadata from the fact-checking articles.
Table 8 shows the prompt instructions provided to GPT4o for the post-processing steps.
@@ Appendix A Labeling prompts
Appendix B Data collection and labeling details
@@
Following the approach of Tonglet et al. (2024), we retrieved all URLs archived on the Wayback Machine777archive.org/help/wayback_api.php under each organization’s web domain. At the time of data collection, all included organizations were current members of the IFCN or under review for membership renewal. We retained URLs containing keywords such as photo, image, or picture, translated into the relevant language. From the raw HTML files, we extracted the article title, publication date, text content, and the URL of the fact-checked image using custom Python scripts. All image URLs were manually verified and corrected when necessary. We manually removed false positives (e.g., non-multimodal claims or fact-checking tutorials), articles containing fact-checking annotations on the image (e.g. an image overlaid with a red cross by the fact-checkers to indicate that it is fake), and duplicate image-claim pairs accross organizations. The final dataset includes 4,982 articles.
GPT4o labeling and post-processing
All GPT4o outputs are generated in English and follow a predefined dictionary format. Task labels and metadata that are not available in the article are annotated as “not enough information”. Additional post-processing steps with GPT4o include: (1) normalizing all dates to the YYYY-MM-DD format, (2) rewriting claims to remove any language revealing the verdict, and (3) extracting the claim in its original language in the article and identifying that language. Instances are taken into account for a given task if GPT4o successfully extracts the corresponding labels. All verdicts are mapped to the following categories: true, false, mixture, unverified. The latter two, which account for only 27 instances, are excluded from verdict prediction. False claims are further categorized into: false (out-of-context), false (manipulated/fake), [^8] Excluding AI-generated images. and false (AI-generated).
@@
For each authentic image labeled as false (out-of-context), we generate a synthetic true claim using GPT4o, combining the following context items, if at least one is available: date, location, people, and event. In the balanced verdict setting, we replace the original false claim with the corresponding synthetic true claim in half of the instances where a true claim was successfully generated. The prompt is provided in Appendix A.
Selecting images for location verification
Images suitable for location verification are identified semi-automatically. First, we use keyword-based filtering to find articles referencing maps, satellite imagery, or detailed geographic verification (e.g., at the street or landmark level). We then manually verify that these images show sufficient background detail to be compared with maps or satellite views, yielding a final set of 65 images. Because the dataset includes only accurate location labels by default, we collect two inaccurate candidates per instance. GPT4o generates a list of likely locations for the image, and the first two with valid GeoNames999geonames.org coordinates are selected as inaccurate candidate locations.
@@
Appendix C Data validation - Prolific study
We recruited annotators on the platform Prolific to validate the quality of the data extracted by GPT4o from fact-checking articles. Annotators were paid at the default rate of £9 per hour. We applied the following criteria to select two annotators per study: they need to (1) be fluent speaker of English and their primary language needs to be the one used by the fact-checking organization, (2) have no language-related disorders, (3) have an approval rate of 99% and (4) have already participated in 50 annotation tasks in the past.
C.1 Annotation guidelines
@@
C.2 Percentage agreement and IAA scores
Table 9 provides the detailed percentages of correct, incorrect, and missing labels, and the IAA scores computed with Randolph’s $\kappa$. Results are provided per type of label and averaged over all 23 Prolific studies. These results exclude the attention tests. For the percentage of correct, incorrect, and missing labels, we report both the lower bound, where a label is considered correct only if both annotators agree on it, and the upper bound, where an answer is considered correct if at least one annotator agrees on it. IAA is reported both with and without taking into account the cases where the label is “not enough information”.
@@ C.2 Percentage agreement and IAA scores
Appendix D Additional dataset statistics
@@
Figure 12 compares the geographic distribution of images with location context items in M4FC and two previous datasets with known location labels, VERITE and 5Pils. VERITE is skewed toward Western contexts, while 5Pils is largely concentrated in Asia and Africa. In contrast, M4FC shows the highest geographic diversity, featuring images from a broad range of cultural contexts. Figure 13 further highlights the geographic diversity of M4FC. However, some areas remain underrepresented, such as East Asia, South America, and Oceania.
@@ 6. Motivation: Why was the image taken?
Figures 14 and 15 show the geographic distribution by country of the claimed and real location of M4FC images, respectively. One map is drawn for each fact-checking organization in M4FC. The maps show that fact-checking organizations tend to focus on images claimed to be associated with their own country. For example, Factly.in and Dfrac.org, both based in India, focus on images with claimed location India. Notable exceptions include Fatabyyano.net, Knack.be, and Dogrula.org. In addition to their own country, many organizations verify images with Ukraine as a claimed location, which reflects the ongoing conflict that started in 2022. For some organizations, the real locations at the country level do not differ significantly from the claimed locations. This is the case for organizations like Factly.in and Indiatoday.in, where all claimed and real locations are within India. This could either mean that location is less frequently misrepresented in out-of-context claims in India, or that the difference in location is at a more fine-grained level, e.g., different cities.
@@
Appendix E Prompt templates for the M4FC tasks
@@
Appendix F Implementation details
@@
Fake detectors fine-tuning
We fine-tune all models for fake detection with the Adam optimizer and a learning rate of 2e52e-5 for 10 epochs.
@@
Table 13: Fake detection results (%). [Uncaptioned image] and [Uncaptioned image] indicate predictions of the base pre-trained model and of the model fine-tuned on M4FC train set, respectively.
@@
For each image, we retrieve up to 20 web pages with the Google RIS engine. We extract their textual content using Trafilatura (Barbaresi, 2021).101010trafilatura.readthedocs.io To prevent evidence leakage, we exclude any web pages published by a curated list of fact-checking organizations, as well as those published after the date of the fact-checking article from which the claim was sourced. For tasks that require text evidence, we treat each web page retrieved with RIS as a single evidence. All 20 text evidence are provided in the prompt, sorted by cosine similarity with the image using CLIP embeddings (Radford et al., 2021).
Appendix G Additional Results
Tables 11,12, 13, and 14 show the results tables on M4FC test set for Visual claim extraction, claimant intent prediction, fake detection, and verdict prediction in the multilingual balanced setting. These tables correspond to the results in Figures 4 and 7.
@@ Appendix G Additional Results
Appendix H Location verification at different zoom levels
@@
Appendix I Pipeline case studies
Figure 17 shows an example of an automated pipeline experiment where the tasks are performed sequentially, using the output of intermediate tasks as input for subsequent ones. The false (out-of-context) claim is the same as in Figure 1. The pipeline starts with claim extraction, which provides a claim that reflects the original social media post well. The intent of the claimant is also accurately interpreted. Based on the social media post, it appears that the claimant is expressing support for Donald Trump. In parallel, the image is recognized as authentic. The output of the next task, image contextualization, is of lower quality. Many errors are introduced, such as the people involved, the event shown, and the date of the picture. The location is correct, although limited to the country level, preventing the step of location verification. However, despite the errors introduced during image contextualization, it is still possible to provide the correct verdict.
Figure 17: An example of an experiment with intermediate outputs. The instance is the same as in Figure 1. Purple boxes indicate (partially) incorrect answers, while green boxes indicate correct answers.

### v2

@@ 5 Conclusion
We introduce M4FC, a real-world dataset for multimodal AFC. The images in M4FC have a broad geographic coverage, representing diverse cultures and events. The dataset supports 10 claim languages and spans six tasks, including two new ones, broadening the scope of multimodal AFC. Our baseline results show that M4FC is a challenging dataset for MLLMs and SOTA AFC methods. Furthermore, we showed that intermediate tasks have a high impact on verdict prediction performance.
@@
We identify five limitations to M4FC.
First, M4FC does not cover videos and audio. Videos have recently become the most prevalent misinformation modality verified by fact-checkers (Dufour et al., 2024), due to the rise of social media platforms like TikTok. However, obtaining the videos used in real-world multimodal claims is technically and legally challenging. The fact-checking articles usually only include a screenshot, and the videos themselves are often removed from the social media platforms where they appeared. Furthermore, scraping and sharing videos, even for research purposes, is prohibited in many countries.
Second, despite the creation of scraping scripts customized for each fact-checking organization, a lot of manual verification was needed. For example, many image URLs had to be corrected. We also did not cover fact-checking web pages that use JavaScript. This limits the possibilities to scale the dataset. Future work should explore more effective methods for extracting the raw content of fact-checking web pages.
@@
Fourth, the balanced and multilingual verdict prediction settings suffer from some biases, as discussed in Appendices A. For evaluation, these biases have little to no impact. However, a trained verdict prediction classifier might exploit these biases to achieve artificially high performance. Hence, it is only recommended to use the English-only imbalanced setting for model training.
Fifth, the location verification task contains 195 instances, which is much less than the other tasks. Only a subset of real-world multimodal misinformation cases is suitable for this task, i.e., a ground or aerial view of a location, and a precise location needs to have been identified in the fact-checking article. Hence, scaling the dataset for this task with real-world cases is not a viable option. Instead, future works should consider creating larger synthetic datasets, e.g., by repurposing large news image dataset with location information, such as TARA (Fu et al., 2022).
@@ Acknowledgments
Appendix A Verdict prediction settings biases
Imbalanced English-only setting
In this default setting, all claims appear as in the fact-checking articles. This setting reflects the real-world distribution of claims verified by fact-checkers, and can be used both for model training and evaluation.
Balanced English-only setting
The balanced English-only setting replaces several false claims with automatically generated true claims based on the true contextual attributes of the image, namely its date, location, event, and depicted people. As a result, the generated true claims typically contain more named entities than real-world claims verified by fact-checkers. In addition, the generated true claims have an average length of 109 characters, whereas real claims in the dataset have an average length of 77 characters. These systematic linguistic differences introduce biases that could be exploited by trained classifiers as superficial heuristics. For this reason, we do not recommend using this setting for model training.
However, we verify that these biases do not affect model evaluation. Table 7 compares verdict prediction performance of two MLLMs under the balanced and imbalanced English-only settings. The resulting differences remain within 3 percentage points. This suggests that, the balanced setting is not less challenging than the imbalanced one as an evaluation setting.
Web evidence Balanced Imbalanced Delta
Qwen2.5VL-7B 69.1 68.8 0.3
72.1 74.3 2.2
GPT4o-mini 74.9 71.9 3.0
77.8 77.6 0.2
Table 7: English-only verdict prediction performance comparison between the balanced and imbalanced settings on M4FC test set, measured with F1 (%). Delta is the absolute difference in F1 score.
Imbalanced multilingual setting
In the multilingual imbalanced setting, most true claims are contained exclusively within two languages, English and Portuguese, while other languages only contain false claims, like Arabic. This is because many non-English speaking fact-checking organizations never write articles about true claims. As a result, a trained model could use the claim’s language as a heuristic for verdict prediction. For this reason, this setting should not be used for model training.
Balanced multilingual setting
In the multilingual balanced setting, both true and false claims are represented for each of the 10 languages. However, the stylistic biases of the synthetic true claims still occurs. For this reason, this setting should not be used for model training.
Appendix B Labeling prompts
Table 8 presents the prompt instructions provided to GPT4o to extract task labels and metadata from the fact-checking articles. Table 9 shows the prompt instructions provided to GPT4o for the post-processing steps.
@@ Appendix B Labeling prompts
Appendix C Data collection and labeling details
@@
Following the approach of Tonglet et al. (2024), we retrieved all URLs archived on the Wayback Machine555archive.org/help/wayback_api.php under each organization’s web domain. At the time of data collection, all included organizations were current members of the IFCN or under review for membership renewal. We retained URLs containing keywords such as photo, image, or picture, translated into the relevant language. From the raw HTML files, we extracted the article title, publication date, text content, and the URL of the fact-checked image using custom Python scripts. All image URLs were manually verified and corrected when necessary. We manually removed false positives (e.g., non-multimodal claims or fact-checking tutorials), articles containing fact-checking annotations on the image (e.g. an image overlaid with a red cross by the fact-checkers to indicate that it is fake), and duplicate image-claim pairs accross organizations. The final dataset includes 4,982 articles.
GPT4o extraction and post-processing
All GPT4o extracted labels are generated in English and follow a predefined dictionary format. Task labels and metadata that are not available in the article are annotated as “not enough information”. Additional post-processing steps with GPT4o include: (1) normalizing all dates to the YYYY-MM-DD format, (2) rewriting claims to remove any words revealing the verdict, and (3) extracting the claim in its original language in the article and identifying that language. For the multilingual verdict prediction setting, we consider languages represented with at least 40 claims; otherwise, we default to the English translation of the claim. All verdicts are mapped to the following categories: true, false, mixture, unverified. The latter two, which account for only 27 instances, are excluded from verdict prediction. False claims are further categorized into: false (out-of-context), false (manipulated/fake), [^6] Excluding AI-generated images. and false (AI-generated).
@@
For each authentic image labeled as false (out-of-context), we generate a synthetic true claim using GPT4o, combining the following context items, if at least one is available: date, location, people, and event. In the balanced verdict setting, we replace the original false claim with the corresponding synthetic true claim in half of the instances where a true claim was successfully generated. The prompt is provided in Appendix B.
@@
Appendix D Data validation - Prolific study
We recruited annotators on the Prolific platform to validate the quality of the data extracted by GPT4o from fact-checking articles. Annotators were paid at the default rate of £9 per hour. We applied the following criteria to select two annotators per study: they need to (1) be fluent speaker of English and their primary language needs to be the one used by the fact-checking organization, (2) have no language-related disorders, (3) have an approval rate of 99% and (4) have already participated in 50 annotation tasks in the past.
D.1 Annotation guidelines
@@
D.2 Percentage agreement and IAA scores
Table 10 provides the detailed percentages of correct, incorrect, and missing labels, and the IAA scores computed with Randolph’s $\kappa$. We distinguish between labels marked as incorrect or missing. A label is marked as missing when GPT4o outputs “not enough information”, despite relevant information being available. Results are provided per type of label and averaged over all 23 Prolific studies. These results exclude the attention tests. For the percentage of correct, incorrect, and missing labels, we report both the lower bound, where a label is considered correct only if both annotators agree on it, and the upper bound, where an answer is considered correct if at least one annotator agrees on it. IAA is reported both with and without taking into account the cases where the label is “not enough information” (NEI). The scores indicate a clear moderate to high IAA. The highest IAA with NEI is 0.81 for the verdict label. For comparison, the IAA between human annotators for verdict in the AVerImaTeC dataset (Cao et al., 2025) is 0.74.
@@ D.2 Percentage agreement and IAA scores
Appendix E Additional dataset statistics
@@
Figure 12 compares the geographic distribution of images with location context items in M4FC and two previous datasets with known location labels, VERITE and 5Pils. VERITE is skewed toward Western contexts, while 5Pils is largely concentrated in Asia and Africa. In contrast, M4FC shows the highest geographic diversity, featuring images from a broad range of cultural contexts. Figure 13 further highlights the geographic diversity of M4FC. However, some areas remain underrepresented, such as East Asia, South America, and Oceania. Figure 14 shows the distribution by country of images with known locations. India and the USA are the most represented, which aligns with their large populations and the strong presence of fact-checking organizations.
@@
Figure 14: Top-20 country distribution of images in M4FC with known location. Countries are represented by their ISO 3166-1 alpha-3 codes.
Task # instances # train # dev # test
VCE 223 115 87 21
CIP 124 82 5 37
FD 322 174 31 117
IC 395 215 35 145
LV 12 0 3 9
VP 322 174 31 117
Table 11: Number of instances per task in the human-validated subset.
@@ 6. Motivation: Why was the image taken?
Figures 15 and 16 show the geographic distribution by country of the claimed and real location of M4FC images, respectively. One map is drawn for each fact-checking organization in M4FC. The maps show that fact-checking organizations tend to focus on images claimed to be associated with their own country. For example, Factly.in and Dfrac.org, both based in India, focus on images with claimed location India. Notable exceptions include Fatabyyano.net, Knack.be, and Dogrula.org. In addition to their own country, many organizations verify images with Ukraine as a claimed location, which reflects the ongoing conflict that started in 2022. For some organizations, the real locations at the country level do not differ significantly from the claimed locations. This is the case for organizations like Factly.in and Indiatoday.in, where all claimed and real locations are within India. This could either mean that location is less frequently misrepresented in out-of-context claims in India, or that the difference in location is at a more fine-grained level, e.g., different cities.
@@
Appendix F Human-validated subset
We release a separate subset of M4FC, called M4FC-validated, which includes only the labels approved by both Prolific annotators during the data validation phase. This subset contains 434 instances in total. Of these, 235 come from the train set, while 38 and 161 come from the dev and test sets, respectively. The task distribution of this subset is reported in Table 11.
Appendix G G-Eval prompts
Tables 13 and 14 provide the G-Eval prompts for visual claim extraction and claimant intent prediction.
Task: Prompt for accessing visual claim extraction results.
Prompt: You will be given:
1. A ground-truth normalized claim written by a human.
2. An AI-generated normalized claim.
Both claims are produced from the same underlying image content.
Your task is to evaluate how consistent and faithful the AI-generated normalized claim is compared with the ground-truth version.
Definition:
Claim Normalization refers to the process of transforming the information contained in an image into a concise textual claim that captures the core semantic meaning. A good normalized claim should accurately reflect the image’s actual content, entities, relationships, and actions.
Evaluation Criteria:
Consistency (1-5):
5 – Fully faithful to the ground-truth claim; preserves all core semantic elements (entities, relations, actions) with no distortions or additions.
4 – Mostly faithful; minor omissions or rephrasing but no incorrect or contradictory content.
3 – Partially faithful; preserves some main elements but misses key relationships or adds ambiguous wording.
2 – Low faithfulness; multiple incorrect or missing semantic elements; contains hallucinations or distorts relationships.
1 – Poor; contradicts the main meaning of the ground-truth claim or fails to represent the image content accurately.
Evaluation Steps:
1. Read the ground-truth normalized claim and identify its core semantic units:
- entities (who/what)
- attributes (descriptions)
- actions/events
- relationships
- spatial or contextual cues
2. Read the AI-generated normalized claim.
3. Compare the semantic overlap and accuracy:
- Does the AI version preserve the core entities?
- Does it distort actions or relationships?
- Does it add hallucinated content not in the ground-truth?
- Does it omit essential elements?
4. Assign a Consistency score (1-5) strictly based on the Evaluation Criteria.
Return STRICT JSON with keys:
{"verdict": one of [consistent, partially_consistent, inconsistent], "score": integer 1-5, "explanation": brief reason}. No extra text.
Evaluation Form (scores ONLY):
- Consistency:
Ground Truth Normalized Claim:
<ground-truth-claim-here>
AI-Generated Normalized Claim:
<ai-generated-claim-here>
Respond with JSON only.
Table 13: G-Eval prompt for VCE task.
Task: Prompt for accessing claimant intent prediction results.
Prompt: You will be given:
1. A ground-truth motivation written by a human.
2. An AI-generated motivation.
Your task is to evaluate how well the AI-generated motivation is consistent and faithful to the meaning and reasoning of the ground-truth version.
Evaluation Criteria:
Consistency (1-5):
5 – Fully faithful, no contradictions, and captures all essential reasoning points.
4 – Mostly faithful; minor omissions but no distortions.
3 – Partially faithful; captures some main points but misses or alters key logic.
2 – Low faithfulness; contains distortions, irrelevant reasoning, or major omissions.
1 – Poor; contradicts or misrepresents the ground-truth motivation.
Evaluation Steps:
1. Read the ground-truth motivation and identify its key reasoning points.
2. Read the AI-generated motivation.
3. Compare meaning, reasoning steps, and any contradictions or hallucinations.
4. Assign a Consistency score from 1 to 5 based on the Evaluation Criteria.
Return STRICT JSON with keys:
{"verdict": one of [consistent, partially_consistent, inconsistent], "score": integer 1-5, "explanation": brief reason}. No extra text.
Evaluation Form (scores ONLY):
- Consistency:
Ground Truth Motivation:
<ground-truth-text-here>
AI-Generated Motivation:
<ai-generated-text-here>
Respond with JSON only.
Table 14: G-Eval prompt for CIP task.
Appendix H Prompt templates for the M4FC tasks
@@
Appendix I Implementation details
@@
Fake image detectors fine-tuning
We fine-tune all models with the Adam optimizer and a learning rate of 2e52e-5 for 10 epochs.
@@
Table 17: Fake image detection results (%). [Uncaptioned image] and [Uncaptioned image] indicate predictions of the base pre-trained model and of the model fine-tuned on M4FC train set, respectively.
@@
For each image, we retrieve up to 20 web pages with the Google RIS engine. We extract their textual content using Trafilatura (Barbaresi, 2021).777trafilatura.readthedocs.io To prevent evidence leakage, we exclude any web pages published by a curated list of fact-checking organizations, as well as those published after the date of the fact-checking article from which the claim was sourced. For tasks that require text evidence, we treat each web page retrieved with RIS as a single evidence. All 20 text evidence are provided in the prompt, sorted by cosine similarity with the image using CLIP embeddings (Radford et al., 2021).
Appendix J Additional Results
Tables 15,16, 17, and 18 show the results tables on M4FC test corresponding to Figures 4 and 7.
@@
Appendix K Location verification at different zoom levels